Betty has a dog-walking service. Last year, she charged $5.00 per dog for each 1/2 mile. This year, she increases the charge per dog by 5%. How much does she earn this year to walk a dog for 1 1/2 miles?
Answer by stuylin(8)

The charge this year = the charge last year + the charge last year * 5 %
The charge this year = 5 + 5 * 0.05
The charge this year = 5 + 0.25
The charge this year = $ 5.25 per 1/2 mile
Number of "1/2 mile" in 1 1/2 miles:
(1 1/2) / (1/2) = 3 "1/2 mile"s
3 * $ 5.25 = $15.75